A spherical asteroid with radius of 18.6 km has uniform density 3500 kg/m3.

Find the gravitational acceleration at the asteroids surface

Express answer in two significant figures
Answer #1

Spherical astroid radius R = 18.6 km = 18600 m
          density d = 3500 kg/m3


   now mass of astroid is m = d*volume of the astroid


  
as astroid is spherical so volume of sphere v = 4/3 pi r^3

                       = 4/3 *3.14*(18.6*10^3)^3
                        = 2.694059712*10^13 m3


mass m = 2.694059712*10^13*3500
   = 9.429208992*10^16 kg


   now gravitational acceleration is g = G*M/R^2


                       g = 6.67408*10^-11 *9.429208992*10^16/(18.6*10^3)^2

                        g = 0.0181903385216 m/s2